OpenStudy (anonymous):

Please help with this Algebra 1 math problem: A system of equations is shown below: 3x - y = 2 x + y = 6 The x-coordinate of the solution to this system of equations is _____.

OpenStudy (texaschic101):

x + y = 6 y = -x + 6 now sub -x + 6 in for y in the other equation 3x - y = 2 3x - (-x + 6) = 2 3x + x - 6 = 2 4x = 2 + 6 4x = 8 x = 2 I am going to find y, just so I can check my answers.. y = -x + 6 y = -2 + 6 y = 4 check.. y = -x + 6 x + y = 6 4 = -2 + 6 4 + 2 = 6 4 = 4(correct) 6 = 6 (correct) solution : x = 2

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Wow thank you for explaining it to me, i get it now haha

OpenStudy (anonymous):

Thank you so much!

OpenStudy (texaschic101):

cool :) glad to help

OpenStudy (anonymous):

Do you think you can help me with another one please?

OpenStudy (anonymous):

Which of the following statements best describes the graph of -5x + 2y = 1? It is a curve joining the points (-5, 2), (2, 3), and (4, 1). It is a curve joining the points (-1, -3), (-1, -3), and (1, 5). It is a straight line joining the points (1, 3), (3, 8), and (-3, -7). It is a straight line joining the points (4, -3), (-1, 2), and (-4, 5).

OpenStudy (texaschic101):

-5x + 2y = 1 Its not A or B... (1,3) x = 1 and y = 3 -5(1) + 2(3) = 1 -5 + 6 = 1 1= 1 (correct) (3,8) x = 3 and y = 8 -5(3) + 2(8) = 1 -15 + 16 = 1 1 = 1 (correct) (-3,-7) x = -3 and y = -7 -5(-3) + 2(-7) = 1 15 - 14 = 1 1 = 1 (correct) answer is C. It is a straight line containing points (1,3) , (3,8) , and (-3,-7) as you can see, I plugged in the points into the equation and they all came out correct

OpenStudy (anonymous):

Oh that makes sense, lol thank you so much :)

OpenStudy (texaschic101):

sure thing :)